# The Composition and Chemistry of Titan’s Atmosphere 111Keywords: Titan; Astrochemistry; Astrobiology; Atmospheres; Photochemistry

### v1

@@ 4.1.11 Diacetylene
The presence of diacetylene (C4H2, butadiyne) was inferred from infrared spectroscopy of Titan’s atmosphere with Voyager’s IRIS instrument 342 via emission bands at 220 and 628 cm-1. At present, it remains the only C4 hydrocarbon species confirmed in Titan’s atmosphere (although note that the nitrile CH3C3N, detected with ALMA168, also has four carbon atoms.)
@@
Future directions: To date, the triacetylene molecule () has remained elusive on Titan, despite its detection in space at high relative abundances compared to C4H2 348. Detection of triacetylene would help to clarify the importance of the radical, which contributes to the depletion of methane in photochemical models along with the smaller related radicals and , as well as the efficacy of polyyne formation in general.
@@ 4.1.12 Benzene
Benzene (c-C6H6, Fig. 21)349, 102, 350, 351, 352 was the second new species detection on Titan made by the Infrared Space Observatory (ISO) in 2003 49, via its strong hydrogen bending mode at 674 cm-1. Benzene was the first cyclic (closed ring) molecule to be detected on Titan, and remains the only confirmed aromatic molecule (molecules with a de-localized electron orbital). The detection of benzene is highly significant, since it provides a measurement of the basic six-membered ring from which larger, multi-ring molecules can be formed,353, 354 building towards macromolecular haze particles - see discussion later in this article.
@@
A second source roughly equal in importance is thought to be dissociative recombination of the C7H ion (benzylium or tropylium, Fig. 22):102
@@
Yet a third ion channel is the DR of C8H with leading to benzene plus other hydrocarbon fragments.102
Radical chemistry also leads to to benzene, such as C2 addition to 1,3 butadiene proposed to occur in the ISM:356
@@
also leads to benzene, but in a highly excited state where it will mostly dissociate to C6H5 + H.102 Recently, a new pathway via a smaller, five-membered ring radical (cyclopentadienyl Fig. 22) molecule has been proposed by Kaiser et al.354:
@@
Loss: Benzene is lost through ionization to the phenylium ion (C6H) and through photolysis to form phenyl ()92, 357. The phenyl radical then either reforms benzene:94
@@ 4.2 Nitrogen Compounds
Nitrogen compounds are formed by chemical combination of dissociation products from initial N2 and CH4, and have formulas . All of the eight known heteroatomic nitrogen compounds are cyanides, wherein nitrogen is bonded to carbon by a triple bond (-CN) and therefore have a formula: -. These are HCN, HNC, CH3CN, C2H3CN, C2H5CN, HC3N, CH3C3N and C2N2. Other than the light molecules HCN and HNC, the remaining molecules are nitriles (organic cyanides). See Fig. 23.
@@ 4.2.1 Nitrogen
Production: The origin of nitrogen in Titan’s atmosphere has been long debated, and is not the subject of this paper. In brief, two major theories exist: enclathratization of N2 gas in the protosolar nebula368, or accretion in the form of NH3 ice followed by later photodissociation to eventually form N2 through a reaction cascade 369:
The latter scenario is currently favored due to the low temperatures in the sub-nebula required to capture molecular nitrogen directly. Variations on the theory include impact conversion of either NH3 or ammonium sulfate () to N2 370, 371.
@@
Isotopes: Since and have significantly different UV cross-sections 377, it is important to correctly account for both isotopes and the wavelength variation of the solar spectrum to arrive at correct dissociation rates. Self-shielding by the more abundant is thought to reduce photolysis rates relative to the less abundant, less shielded , causing a lower ratio in nitrogen atoms than in the original molecules. Since significant amounts of atomic nitrogen go on to form nitriles, this skew towards increased production of may explain the lower in nitriles than in N2 itself 377, 378, 379.
Future directions: The dissociation and reaction pathways for N2 and its daughter ions and radicals remain one of the better known areas of Titan chemistry. However gaps remain, in particular whether nitrogen exists in chemicals such as amines and imines, or if it is incorporated in to heterocyclic ring molecules. This is further discussed in a later section.
@@ 4.2.2 Hydrogen Cyanide
Hydrogen cyanide was first detected by the Voyager 1 IRIS spectrometer through its strong infrared emission at 712 cm-1 380, and later at sub-millimeter wavelengths from ground-based observatories 381, 382, 383, 164. Although a relatively simple molecule that has been included in photochemical models for more than four decades, gaps in our knowledge of HCN formation may still exist, and new pathways have been identified recently 384.
@@
HCN has been shown to form co-crystals with hydrocarbons at Titan-relevant temperatures 389, the study of which will be important for understanding the solids and liquids on the surface. HCN, along with HC3N, has also been implicated in the formation of C4N2 in grain-surface chemical reactions 390, which requires further study to elucidate reaction rates and whether this process is sufficient to explain observed ice spectral properties 391.
Finally, HCN has been implicated in processes of astrobiological importance. A well known example is its proposed ability to directly form the amino acid adenine (C5H5N5)392, 393, 394, 395 from the rearrangement (oligomerization) of five HCN molecules. Although the importance of this reaction for the seeding of life on the early Earth has been disputed396, 397, it may be more prevalent on Titan where HCN occurs in greater abundance.398, 399 HCN may also have the potential to polymerize into polyimines, structures that may catalyze astrobiologically important reactions 400. The astrobiological potential of HCN therefore remains under continued investigation401, 384, 402, 403.
@@ 4.2.3 Hydrogen Isocyanide
with R188 dominating in the thermosphere and R189 becoming important in the mesosphere and below.90 Ion pathways may also be similar (see Fig. 26)385, although branching ratios are in most cases more uncertain than for HCN, e.g. through dissociative recombination of HCNH+:407
HNC may also be produced as photodissociation product of C2H3CN387 in the upper atmosphere, and a further production peak may occur due to cosmic ray chemistry at 100-150 km 98.
@@
Future directions: HNC/HCN is now one of two isomer pairs known in Titan’s atmosphere (the other being ). Study of the branching ratios and reaction rates leading to and from isomer pairs/triples etc is of importance because the less stable isomer(s) may follow different reaction pathways compared to the more abundant molecule(s). Therefore for a complete understanding of Titan’s atmospheric chemistry, all isomers must be included in models. Study of the vertical ratio between HCN/HNC and CH3CCH/CH2CCH2 may also provide useful information on the abundance of atomic H, collisions which can cause conversion between the isomers.
@@ 4.2.4 Acetonitrile
Acetonitrile was first detected on Titan in the early 1990s by millimeter wavelength astronomy 408, followed ten years later by the first measurement of its vertical profile 164 using the 30 m telescope at IRAM. CH3CN was the first Titan molecule to be first detected at millimeter wavelengths, an astronomical technique that was to yield many other discoveries later with ALMA.
@@
and by the termolecular reaction of H with cyanomethyl (CH2CN), in a chain that begins with acrylonitrile (C2H5CN):98, 103
@@ 4.2.5 Cyanoacetylene
While HC3N does react without a barrier with radicals such as CN and C2H, the main loss channel in the neutral atmosphere is thought to be successive hydrogen addition:98
@@ 4.2.6 Cyanogen
Cyanogen (C2N2), like cyanoacetylene was first detected in Titan’s atmosphere by the Voyager IRIS spectrometer in the infrared 342 at 233 cm-1.
@@ 4.2.7 Acrylonitrile
Acrylonitrile (C2H3CN) was the second molecule to be discovered on Titan at millimeter wavelengths using ALMA 166, following the detection of propionitrile,167 discussed in the next section.
@@
Loss: In a similar manner to HCN, and other nitriles, C2H3CN is lost in the ionosphere by the two-step process of proton transfer:433, 434
@@ 4.2.8 Propionitrile
Propionitrile (C2H5CN) was the second molecule to be originally detected using sub-millimeter astronomy and the first molecule with ALMA 167.
@@
Future directions: Propionitrile has been asserted to condense in pure crystalline form in Titan’s atmosphere 441, on the basis of an unexplained feature in Titan’s far-infrared spectrum. This has been questioned based on vapor pressure of the gaseous form 442, 443, although it is possible that a co-condensed ice containing C2H5CN along with other gases may replicate the unexplained ‘haystack’ emission 444. Further work on spectroscopy will be required to determine if this is a unique solution, or if other possibilities exist.
@@ 4.2.9 Cyanopropyne
Cyanopropyne CH3C3N was the fourth molecule to be discovered by ALMA spectroscopy of Titan at mm wavelengths 168, following previous detection in the ISM 445.
@@
Future directions: , has at least three stable isomers that have been detected in space 450, 451. Besides the currently detected isomer (cyanopropyne, CH3C3N, butynenitrile or methylcyanoacetylene), there is also cyanoallene ()452 and propargyl cyanide ()453, both being first detected in the Taurus Molecular Cloud (TMC-1) at radio wavelengths. These provide good targets for detection on Titan, and their measurement would help to constrain photochemical pathways and models. Further more exotic arrangements of the same atoms may also exist and remain to be detected.
@@ 4.3.1 Water
while in the middle atmosphere it reacts with CO to form CO2: 97
@@ 4.3.2 Carbon Monoxide
Recently, high-sensitivity observations with ALMA have narrowed the experimental error to a range of ppm 133, making it the fourth most abundance species in Titan’s atmosphere after N2, CH4 and H2. As a triple-bonded molecule, CO once produced is both resistant to photolysis and chemical reaction, as well having no loss through condensation, and hence all evidence available at present points to a uniform vertical profile.
@@
(Formaldehyde may also be created by reaction of OH with 3CH2, C2H4 etc).
(also OH + 3CH2, OH + C2H4 etc). Note that formaldehyde has yet to be detected in Titan’s atmosphere.
@@
Loss: CO is primarily lost slowly through reaction with OH forming CO2:474
@@ 4.3.3 Carbon Dioxide
Carbon dioxide (CO2) was one of seven new gas species detected by Voyager’s IRIS infrared spectrometer 476, 477 and subsequently by Cassini CIRS 478, 278. Unlike shorter-lived chemical chemical species ( HC3N, C4H2), CO2 exhibits little variation with latitude in the lower stratosphere, lacking a polar enhancement.
Production: CO2 is thought to be mostly produced from CO + OH as shown in the previous section (R 255). See also Fig. 36464, 479, 103.
@@ 4.4 Growth of large particles
HACA - "hydrogen abstraction- C2H2 addition"490
@@
Once H- is produced, it leads to the creation of some larger negative ions through proton abstraction:103
@@ 5.1.1 Hydrocarbons
At present, C4H2 remains the only hydrocarbon species definitively detected in the neutral atmosphere, while no aliphatic species with have been detected (benzene, c-C6H6, being detected as a ring).
@@
Upper limits for -butane and -butane (C4H10) have been calculated as and respectively, from CIRS spectra at 200–250 km, 30∘N–30∘S498.
@@
In parallel, photochemical models that currently treat multiple isomers under single formulae such as and must continue to expand treatment of separate isomers. To date this has been sparse and primarily for a few specific cases for which isomeric data exists: especially HCN and HNC, and CH2CCH2 and CH3CCH, which are considered separately in current models 90, 91, 99, 103. A good reason for lack of inclusion of separate isomers is lack of knowledge of branching ratios and isomer-specific reaction rates. Monte Carlo simulations have proved useful at showing which reaction rate uncertainties have the biggest effects on the uncertainties in the solution 87, 88. These researches thereby provide important prioritization of where resources such as laboratory time and theoretical effort (e.g. TST calculations) can best be spent to most rapidly improve our knowledge.
@@ 5.1.2 Nitrogen Species
All of the nine detected nitrogen compounds in Titan’s neutral atmosphere, formed from dissociation products of N2 and CH4, are triple bonded. These include N2 itself, and eight known cyanides, which include nitrogen in a terminal -CN functional group.
@@
Nitrogen has yet to be detected in other types of bonding - i.e. where it is not in a triple bond, such as in amines, imines (beyond HNC), azines and nitrogen heterocycles (see Fig. 40). A simple example is ammonia (NH3) which was tentatively inferred from Cassini’s mass spectra at high altitudes, at mixing fractions of , although not yet uniquely detected due to the barometric degeneracy problem of unit resolution mass spectroscopy.501, 219
The major channel for ammonia production is thought to be via NH2 + CH2N, as follows:103
@@
Upper limits for some molecules in the lower atmosphere have been estimated, including for ammonia (NH3) of 1.3 ppb (3-) at 107 km, 25∘S 502, and methanimine (CH2NH) at 0.35 ppb (3-) in the stratosphere 503.
Future Work: No molecules having both oxygen and nitrogen, e.g. HNO, and more complex amino acids, have been detected. Detection of such functional groups and molecules, important for biological activity on Earth, is a key area of future research. Several nitriles have been detected only through sub-millimeter astronomical techniques as previously mentioned, but not using infrared techniques despite several decades of attempts 499, 500, 49, 502. In fact laboratory line lists with intensities do not exist for many of the nitriles sought in the infrared, except CH3CN 504, 505, hampering the search and implying a need for new laboratory work to obtain cold temperature spectra.

### v2

@@ 4.1.11 Diacetylene
The presence of diacetylene (CH, butadiyne) was inferred from infrared spectroscopy of Titan’s atmosphere with Voyager’s IRIS instrument 342 via emission bands at 220 and 628 cm. At present, it remains the only C hydrocarbon species confirmed in Titan’s atmosphere (although note that the nitrile CHCN, detected with ALMA168, also has four carbon atoms.)
@@
Future directions: To date, the triacetylene molecule () has remained elusive on Titan, despite its detection in space at high relative abundances compared to CH 348. Detection of triacetylene would help to clarify the importance of the radical, which contributes to the depletion of methane in photochemical models along with the smaller related radicals and , as well as the efficacy of polyyne formation in general.
@@ 4.1.12 Benzene
Benzene (c-CH, Fig. 21)349, 102, 350, 351, 352 was the second new species detection on Titan made by the Infrared Space Observatory (ISO) in 2003 49, via its strong hydrogen bending mode at 674 cm. Benzene was the first cyclic (closed ring) molecule to be detected on Titan, and remains the only confirmed aromatic molecule (molecules with a de-localized electron orbital). The detection of benzene is highly significant, since it provides a measurement of the basic six-membered ring from which larger, multi-ring molecules can be formed,353, 354 building towards macromolecular haze particles - see discussion later in this article.
@@
A second source roughly equal in importance is thought to be dissociative recombination of the CH ion (benzylium or tropylium, Fig. 22):102
@@
Yet a third ion channel is the DR of CH with leading to benzene plus other hydrocarbon fragments.102
Radical chemistry also leads to to benzene, such as C addition to 1,3 butadiene proposed to occur in the ISM:356
@@
also leads to benzene, but in a highly excited state where it will mostly dissociate to CH + H.102 Recently, a new pathway via a smaller, five-membered ring radical (cyclopentadienyl Fig. 22) molecule has been proposed by Kaiser et al.354:
@@
Loss: Benzene is lost through ionization to the phenylium ion (CH) and through photolysis to form phenyl ()92, 357. The phenyl radical then either reforms benzene:94
@@ 4.2 Nitrogen Compounds
Nitrogen compounds are formed by chemical combination of dissociation products from initial N and CH, and have formulas . All of the eight known heteroatomic nitrogen compounds are cyanides, wherein nitrogen is bonded to carbon by a triple bond (-CN) and therefore have a formula: -. These are HCN, HNC, CHCN, CHCN, CHCN, HCN, CHCN and CN. Other than the light molecules HCN and HNC, the remaining molecules are nitriles (organic cyanides). See Fig. 23.
@@ 4.2.1 Nitrogen
Production: The origin of nitrogen in Titan’s atmosphere has been long debated, and is not the subject of this paper. In brief, two major theories exist: enclathratization of N gas in the protosolar nebula368, or accretion in the form of NH ice followed by later photodissociation to eventually form N through a reaction cascade 369:
The latter scenario is currently favored due to the low temperatures in the sub-nebula required to capture molecular nitrogen directly. Variations on the theory include impact conversion of either NH or ammonium sulfate () to N 370, 371.
@@
Isotopes: Since and have significantly different UV cross-sections 377, it is important to correctly account for both isotopes and the wavelength variation of the solar spectrum to arrive at correct dissociation rates. Self-shielding by the more abundant is thought to reduce photolysis rates relative to the less abundant, less shielded , causing a lower ratio in nitrogen atoms than in the original molecules. Since significant amounts of atomic nitrogen go on to form nitriles, this skew towards increased production of may explain the lower in nitriles than in N itself 377, 378, 379.
Future directions: The dissociation and reaction pathways for N and its daughter ions and radicals remain one of the better known areas of Titan chemistry. However gaps remain, in particular whether nitrogen exists in chemicals such as amines and imines, or if it is incorporated in to heterocyclic ring molecules. This is further discussed in a later section.
@@ 4.2.2 Hydrogen Cyanide
Hydrogen cyanide was first detected by the Voyager 1 IRIS spectrometer through its strong infrared emission at 712 cm 380, and later at sub-millimeter wavelengths from ground-based observatories 381, 382, 383, 164. Although a relatively simple molecule that has been included in photochemical models for more than four decades, gaps in our knowledge of HCN formation may still exist, and new pathways have been identified recently 384.
@@
HCN has been shown to form co-crystals with hydrocarbons at Titan-relevant temperatures 389, the study of which will be important for understanding the solids and liquids on the surface. HCN, along with HCN, has also been implicated in the formation of CN in grain-surface chemical reactions 390, which requires further study to elucidate reaction rates and whether this process is sufficient to explain observed ice spectral properties 391.
Finally, HCN has been implicated in processes of astrobiological importance. A well known example is its proposed ability to directly form the amino acid adenine (CHN)392, 393, 394, 395 from the rearrangement (oligomerization) of five HCN molecules. Although the importance of this reaction for the seeding of life on the early Earth has been disputed396, 397, it may be more prevalent on Titan where HCN occurs in greater abundance.398, 399 HCN may also have the potential to polymerize into polyimines, structures that may catalyze astrobiologically important reactions 400. The astrobiological potential of HCN therefore remains under continued investigation401, 384, 402, 403.
@@ 4.2.3 Hydrogen Isocyanide
with R188 dominating in the thermosphere and R189 becoming important in the mesosphere and below.90 Ion pathways may also be similar (see Fig. 26)385, although branching ratios are in most cases more uncertain than for HCN, e.g. through dissociative recombination of HCNH:407
HNC may also be produced as photodissociation product of CHCN387 in the upper atmosphere, and a further production peak may occur due to cosmic ray chemistry at 100-150 km 98.
@@
Future directions: HNC/HCN is now one of two isomer pairs known in Titan’s atmosphere (the other being ). Study of the branching ratios and reaction rates leading to and from isomer pairs/triples etc is of importance because the less stable isomer(s) may follow different reaction pathways compared to the more abundant molecule(s). Therefore for a complete understanding of Titan’s atmospheric chemistry, all isomers must be included in models. Study of the vertical ratio between HCN/HNC and CHCCH/CHCCH may also provide useful information on the abundance of atomic H, collisions which can cause conversion between the isomers.
@@ 4.2.4 Acetonitrile
Acetonitrile was first detected on Titan in the early 1990s by millimeter wavelength astronomy 408, followed ten years later by the first measurement of its vertical profile 164 using the 30 m telescope at IRAM. CHCN was the first Titan molecule to be first detected at millimeter wavelengths, an astronomical technique that was to yield many other discoveries later with ALMA.
@@
and by the termolecular reaction of H with cyanomethyl (CHCN), in a chain that begins with acrylonitrile (CHCN):98, 103
@@ 4.2.5 Cyanoacetylene
While HCN does react without a barrier with radicals such as CN and CH, the main loss channel in the neutral atmosphere is thought to be successive hydrogen addition:98
@@ 4.2.6 Cyanogen
Cyanogen (CN), like cyanoacetylene was first detected in Titan’s atmosphere by the Voyager IRIS spectrometer in the infrared 342 at 233 cm.
@@ 4.2.7 Acrylonitrile
Acrylonitrile (CHCN) was the second molecule to be discovered on Titan at millimeter wavelengths using ALMA 166, following the detection of propionitrile,167 discussed in the next section.
@@
Loss: In a similar manner to HCN, and other nitriles, CHCN is lost in the ionosphere by the two-step process of proton transfer:433, 434
@@ 4.2.8 Propionitrile
Propionitrile (CHCN) was the second molecule to be originally detected using sub-millimeter astronomy and the first molecule with ALMA 167.
@@
Future directions: Propionitrile has been asserted to condense in pure crystalline form in Titan’s atmosphere 441, on the basis of an unexplained feature in Titan’s far-infrared spectrum. This has been questioned based on vapor pressure of the gaseous form 442, 443, although it is possible that a co-condensed ice containing CHCN along with other gases may replicate the unexplained ‘haystack’ emission 444. Further work on spectroscopy will be required to determine if this is a unique solution, or if other possibilities exist.
@@ 4.2.9 Cyanopropyne
Cyanopropyne CHCN was the fourth molecule to be discovered by ALMA spectroscopy of Titan at mm wavelengths 168, following previous detection in the ISM 445.
@@
Future directions: , has at least three stable isomers that have been detected in space 450, 451. Besides the currently detected isomer (cyanopropyne, CHCN, butynenitrile or methylcyanoacetylene), there is also cyanoallene ()452 and propargyl cyanide ()453, both being first detected in the Taurus Molecular Cloud (TMC-1) at radio wavelengths. These provide good targets for detection on Titan, and their measurement would help to constrain photochemical pathways and models. Further more exotic arrangements of the same atoms may also exist and remain to be detected.
@@ 4.3.1 Water
while in the middle atmosphere it reacts with CO to form CO: 97
@@ 4.3.2 Carbon Monoxide
Recently, high-sensitivity observations with ALMA have narrowed the experimental error to a range of ppm 133, making it the fourth most abundance species in Titan’s atmosphere after N, CH and H. As a triple-bonded molecule, CO once produced is both resistant to photolysis and chemical reaction, as well having no loss through condensation, and hence all evidence available at present points to a uniform vertical profile.
@@
(Formaldehyde may also be created by reaction of OH with CH, CH etc).
(also OH + CH, OH + CH etc). Note that formaldehyde has yet to be detected in Titan’s atmosphere.
@@
Loss: CO is primarily lost slowly through reaction with OH forming CO:474
@@ 4.3.3 Carbon Dioxide
Carbon dioxide (CO) was one of seven new gas species detected by Voyager’s IRIS infrared spectrometer 476, 477 and subsequently by Cassini CIRS 478, 278. Unlike shorter-lived chemical chemical species ( HCN, CH), CO exhibits little variation with latitude in the lower stratosphere, lacking a polar enhancement.
Production: CO is thought to be mostly produced from CO + OH as shown in the previous section (R 255). See also Fig. 36464, 479, 103.
@@ 4.4 Growth of large particles
HACA - "hydrogen abstraction- CH addition"490
@@
Once H is produced, it leads to the creation of some larger negative ions through proton abstraction:103
@@ 5.1.1 Hydrocarbons
At present, CH remains the only hydrocarbon species definitively detected in the neutral atmosphere, while no aliphatic species with have been detected (benzene, c-CH, being detected as a ring).
@@
Upper limits for -butane and -butane (CH) have been calculated as and respectively, from CIRS spectra at 200–250 km, 30N–30S498.
@@
In parallel, photochemical models that currently treat multiple isomers under single formulae such as and must continue to expand treatment of separate isomers. To date this has been sparse and primarily for a few specific cases for which isomeric data exists: especially HCN and HNC, and CHCCH and CHCCH, which are considered separately in current models 90, 91, 99, 103. A good reason for lack of inclusion of separate isomers is lack of knowledge of branching ratios and isomer-specific reaction rates. Monte Carlo simulations have proved useful at showing which reaction rate uncertainties have the biggest effects on the uncertainties in the solution 87, 88. These researches thereby provide important prioritization of where resources such as laboratory time and theoretical effort (e.g. TST calculations) can best be spent to most rapidly improve our knowledge.
@@ 5.1.2 Nitrogen Species
All of the nine detected nitrogen compounds in Titan’s neutral atmosphere, formed from dissociation products of N and CH, are triple bonded. These include N itself, and eight known cyanides, which include nitrogen in a terminal -CN functional group.
@@
Nitrogen has yet to be detected in other types of bonding - i.e. where it is not in a triple bond, such as in amines, imines (beyond HNC), azines and nitrogen heterocycles (see Fig. 40). A simple example is ammonia (NH) which was tentatively inferred from Cassini’s mass spectra at high altitudes, at mixing fractions of , although not yet uniquely detected due to the barometric degeneracy problem of unit resolution mass spectroscopy.501, 219
The major channel for ammonia production is thought to be via NH + CHN, as follows:103
@@
Upper limits for some molecules in the lower atmosphere have been estimated, including for ammonia (NH) of 1.3 ppb (3-) at 107 km, 25S 502, and methanimine (CHNH) at 0.35 ppb (3-) in the stratosphere 503.
Future Work: No molecules having both oxygen and nitrogen, e.g. HNO, and more complex amino acids, have been detected. Detection of such functional groups and molecules, important for biological activity on Earth, is a key area of future research. Several nitriles have been detected only through sub-millimeter astronomical techniques as previously mentioned, but not using infrared techniques despite several decades of attempts 499, 500, 49, 502. In fact laboratory line lists with intensities do not exist for many of the nitriles sought in the infrared, except CHCN 504, 505, hampering the search and implying a need for new laboratory work to obtain cold temperature spectra.